Given the functions:
f(x)= {(3,5), (2,4), (1,7)}
h(x)= {(3,2), (4,3), (1,6)}
g(x)= sqrt(x-3)
k(x)= x^2=5
determine
1) f(h(3))
2) g(k(x))
3) g(k(7))
4) k(g(x))

: Given the functions:
f(x)= {(3,5), (2,4), (1,7)}
h(x)= {(3,2), (4,3), (1,6)}
g(x)= sqrt(x-3)
k(x)= x^2=5
determine
1) f(h(3))
2) g(k(x))
3) g(k(7))
4) k(g(x))


Answer by Fombitz(1740)

1) h(3)=2
f(h(3))=f(2)=4
.
.
.
2)k(x)=x^2-5
g(x)= sqrt(x-3)
g(k(x))= sqrt((x^2-5)-3)
g(k(x))= sqrt(x^2-8)
.
.
.
3) From 2, g(k(x))= sqrt(x^2-8)
g(k(7))= sqrt(7^2-8)
g(k(7))= sqrt(49-8)
g(k(7))= sqrt(41)
.
.
.
4)g(x)=sqrt(x-3)
k(x)=x^2-5
k(g(x))=(sqrt(x-3))^2-5
k(g(x))=(x-3)-5
k(g(x))=x-8